OpenStudy (anonymous):

4. Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph.

OpenStudy (anonymous):

@austinL

OpenStudy (campbell_st):

well the easiest way if you know the roots of zeros is find the midpoint between the zeros... this will be the line of symmetry... x = h then substitute the value into the original equation... and get y = k so the vertex will be (h, k) if you have an equation in the form \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\] then substitute it to find y... hope it helps

OpenStudy (anonymous):

But I don't understand how to solve one of those equations.....

OpenStudy (campbell_st):

oops forgot to say, parabolas are symmetric... and the vertex lies on the line of symmetry

OpenStudy (anonymous):

Here's my work: f(x) = 3x2 - x + 5 f(x) = 3(x2 – x) + 5 That's all I could do.

OpenStudy (anonymous):

*3x^2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so you're attempting to put the equation into vertex from I wouldn't worry using the line of symmetry \[x = \frac{-(-1)}{6}\] substitute it to find y = 59/12 vertex (1/6, 59/12) I just find this easier then completing the square

OpenStudy (anonymous):

My problem is how do I solve this? f(x) = 3x^2 - x + 5

OpenStudy (campbell_st):

well you don't the curve is positive definite... it doesn't cut the x axis.. so it has complex roots... you can test by using the discriminant b^2 - 4ac (-1)^2 - 4 *3 * 5 = -59 so the equation have no real roots...